# Value-Guided Search for Efficient Chain-of-Thought Reasoning

### v1

@@ Appendix D Further Details of Data Collection
Roll-in roll-out process. We also provide further intuition for the roll-in vs. roll-out process (illustrated in Fig. 2 left). The roll-in and then roll-out process is a standard technique in imitation learning [12] and reinforcement learning [11].
@@
Roll-in. The main point of the roll-in process is to create a diverse distribution of partial solutions to roll-out from. By creating a diverse roll-in distribution with multiple roll-in policies, we can ensure that the classifier is trained on a diverse context distribution and will generalize better to new traces. To select where to cut a roll-in (to start the roll-out), we sample a cut index from the distribution of $p(i)=\frac{\sqrt{i}}{\sum_{j\in[L]}\sqrt{j}}$ where $L$ is the length of the roll-in. We chose this such that the cut index is more likely to occur at earlier positions of the roll-in. We want to encourage more learning at earlier positions since those prediction problems are more difficult than at later positions. The following figure (Fig. 16) illustrates the distribution of the length of roll-outs, which shows that this indexing scheme indeed yields many long roll-outs.
@@ Appendix E Further Details for Value Model Training
Score Head Labels 00=Incorrect, $1$=Correct, $2$=Incomplete
@@
Grad Norm Clip $5$
[…]
@@ Appendix G Further Details for Training Bradley-Terry Reward Model
Dataset. Recall that our dataset for value model training (OpenR1-VM) contains $56$ responses per problem. To construct a Bradley-Terry dataset, we sample up to $4$ response pairs per problem, where a pair consists of a response with reward 00 (the ‘reject’ response) and a response with reward $1$ (the ‘chosen’ response). Some prompts may have fewer than $4$ responses with reward 00 or $1$, and in those cases, we include as many as possible. This yields a dataset of roughly $122$k pairs.
Model. We use the same model architecture as the value model (Appendix E) except that the score head outputs a single scalar score instead of a three-dimensional vector. We use the same training pipeline, except that the training loss is swapped to the standard BT loss [8]:
@@ Appendix H Inference FLOPS Computation
AIME-24 Problem: Jen enters a lottery by picking $4$ distinct numbers from $S=\{1,2,3,\cdots,9,10\}.$ $4$ numbers are randomly chosen from $S.$ She wins a prize if at least two of her numbers were $2$ of the randomly chosen numbers, and wins the grand prize if all four of her numbers were the randomly chosen numbers. The probability of her winning the grand prize given that she won a prize is $\tfrac{m}{n}$ where $m$ and $n$ are relatively prime positive integers. Find $m+n$. Rejected beam with score 0.009: Okay, so I need to figure out the probability that Jen wins the grand prize given that she won a prize. That means I have to use conditional probability here. I remember that conditional probability is $P(A|B)=P(A\cap B)/P(B)$. In this case, event A is her winning the grand prize, and event B is her winning a prize. So, I need to find the probability that she won both the grand prize and a prize, which is just her winning the grand prize, divided by the probability that she won a prize. First, let me figure out the total number of possible ways Jen can pick 4 numbers from S, which has 10 numbers. That should be the combination of 10 choose 4, right? So, $C(10,4)=210$. So, the total number of possible combinations is 210. So, she can win a prize in two scenarios: either exactly two of her numbers are the randomly chosen ones, or all four are. …The number of ways she can have exactly two numbers correct is $C(4,2)$ ways to choose the correct numbers, and then $C(6,2)$ ways to choose the incorrect ones from the remaining 6 numbers. Similarly, the number of ways she can have all four numbers correct is just $C(4,4)=1$. So, the total number of winning ways is $C(4,2)*C(6,2)+C(4,4)$. …$C(4,2)$ is 6, and $C(6,2)$ is 15. So, 6*15 = 90. Then, $C(4,4)$ is 1. So, total winning ways are 90 + 1 = 91. Alternatively, I can compute the probability of not winning a prize, which is the probability that she has 0 or 1 correct numbers. Let me compute that and subtract from 1 to see if it matches. Number of ways to have 0 correct numbers: $C(4,0)*C(6,4)=1*15=15$. Number of ways to have 1 correct number: $C(4,1)*C(6,3)=4*20=80$. So, total non-winning ways: 15 + 80 = 95. Therefore, probability of not winning a prize is 95/210, so probability of winning a prize is $1-95/210=(210-95)/210=115/210$. Wait, that doesn’t match with the previous result. Hmm, something’s wrong here. Wait, why did I get 91 before? Let me see. I had $C(4,2)*C(6,2)+C(4,4)$. So, that’s 6*15 +1=91. But that’s only the number of winning ways where she has exactly two or exactly four correct numbers. So, 91 winning ways, 95 non-winning, which adds to 186, which is less than 210. So, the rest must be the impossible case? No, because in reality, she can’t have 0 correct numbers if she has 4 numbers. Wait, no, she can have 0 correct numbers if all her numbers are among the 6 incorrect ones. So, that’s 15 ways. Similarly, 1 correct number is 80 ways, and 2 correct numbers is 90 ways, and 4 correct numbers is 1 way. So, 15 +80 +90 +1=186. Wait, so actually, 186 winning ways and 24 non-winning ways. So, probability of winning is 186/210, which simplifies to 31/35. … Therefore, $P(A|B)=(1/210)/(186/210)=1/186$. So, m is 1 and n is 186. Are they relatively prime? 1 and 186 share no common factors besides 1, so yes. So, $m+n=1+186=187$. … Selected Beam with score 0.996: Okay, so this is a conditional probability problem. The formula for conditional probability is $P(A|B)=P(A\cap B)/P(B)$, where A is the event that she wins the grand prize, and B is the event that she wins a prize. Since winning the grand prize is a specific case of winning a prize, A is a subset of B. So, $P(A\cap B)$ is just $P(A)$. Therefore, the probability we’re looking for is $P(A)/P(B)$. … Since she picks 4 numbers from 10, the total number of possible 4-number combinations is $C(10,4)$. The number of favorable outcomes for A is 1 (since there’s only one exact set of four numbers that could be the randomly chosen ones). So, $P(A)$ is $1/C(10,4)$. … To find $P(B)$, it might be easier to compute the probability of the complementary event, which is that Jen wins zero prizes or one prize, and subtract that from 1. So, $P(B)$ = 1 - P(0 prizes) - P(1 prize). Let me compute P(0 prizes) and P(1 prize). First, P(0 prizes): This is the probability that none of Jen’s four numbers are among the four randomly chosen. So, all four numbers she picked are from the remaining six numbers (since S has 10 numbers, and if she picks 4, the remaining are 6). So, the number of ways to choose 4 numbers from 6 is $C(6,4)$. Therefore, P(0 prizes) = $C(6,4)/C(10,4)$. Similarly, P(1 prize): This is the probability that exactly one of Jen’s numbers is among the four randomly chosen. So, we need to choose 1 number from her 4, and 3 numbers from the remaining 6 (since 10 - 4 = 6). So, the number of ways is $C(4,1)*C(6,3)$. Therefore, P(1 prize) = $[C(4,1)*C(6,3)]/C(10,4)$. …So, putting it all together: $P(B)=1-[C(6,4)/C(10,4)]-[C(4,1)*C(6,3)/C(10,4)]$ Therefore, $P(B)=1-15/210-80/210$. …Therefore, $P(B)$ is 23/42. …Therefore, the conditional probability $P(A|B)$ is $(1/210)/(23/42)=(1/210)*(42/23)=(42)/(210*23)$Therefore, the probability becomes $(1/5)/23=1/(5*23)=1/115$. …But 1/115 can’t be simplified further, so m is 1 and n is 115, which are coprime. Therefore, m + n is 1 + 115 = 116. …
@@
AIME-25 Problem: Find the sum of all positive integers $n$ such that $n+2$ divides the product $3(n+3)(n^{2}+9)$. Rejected beam with score 0.395: First, let’s parse the problem. I need to find all positive integers $n$ where $n+2$ is a divisor of $3(n+3)(n^{2}+9)$. So, mathematically, this means that $n+2$ divides $3(n+3)(n^{2}+9)$. In other words, when you divide $3(n+3)(n^{2}+9)$ by $n+2$, there should be no remainder. So, $n+2$ must be a factor of this product. One way to approach this is by using the Factor Theorem. The Factor Theorem states that if $n+2$ is a factor of a polynomial $P(n)$, then $P(-2)=0$. So, if I let $P(n)=3(n+3)(n^{2}+9)$, then $P(-2)$ should be equal to zero if $n+2$ divides $P(n)$. Let me compute $P(-2)$: $P(-2)=3(-2+3)((-2)^{2}+9)$ Simplify each part: $(-2+3)=1$ $(-2)^{2}=4$, so $4+9=13$ Thus, $P(-2)=3*1*13=39$Wait, perhaps I’m misapplying the Factor Theorem here. Let me think again. Alternatively, another approach is to express $n+2$ as $k$, so $k=n+2$. Then, $n=k-2$. Substitute $n=k-2$ into the expression $3(n+3)(n^{2}+9)$ and see what conditions must hold for $k$ to divide this. …Therefore, the product becomes: $3(k+1)(k^{2}-4k+13)$ Which implies that $k$ divides $3(k+1)(k^{2}-4k+13)$. Since $k$ divides the product, it must divide at least one of the factors. But $k$ divides $k$, so we can consider the other factors. But wait, $k$ divides $3(k+1)(k^{2}-4k+13)$, so $k$ divides $3(k+1)(k^{2}-4k+13)$. Since $k$ and $k+1$ are consecutive integers, their GCD is 1. Similarly, $k$ and $k^{2}-4k+13$ may have a GCD. Let me check: Let me denote $d=\gcd(k,k^{2}-4k+13)$. Then $d$ divides $k$ and $k^{2}-4k+13$. So, $d$ divides $k^{2}-4k+13-k(k-4)=k^{2}-4k+13-k^{2}+4k=13$. So, $d$ divides 13. Therefore, $d=1$ or $13$. Wait, maybe I’m complicating things. Let me approach this differently. … But maybe there’s another way to approach this. Let me try plugging in small positive integers for $n$ and see if $n+2$ divides the product $3(n+3)(n^{2}+9)$. That might help confirm if $n=1$ is the only solution. Testing $n=1$: As above, product is 120, which is divisible by 3. Testing $n=2$: $n+2=4$. Product is $3(5)(13)=195$. 195 divided by 4 is 48.75, which isn’t an integer. So, 4 doesn’t divide 195. Testing $n=9$: $n+2=11$. Product is $3(12)(90)=3240$. 3240 divided by 11 is approximately 294.545, not integer. Testing $n=10$: $n+2=12$. Product is $3(13)(109)=4209$. 4209 divided by 12 is 350.75, not integer. Wait, so up to $n=10$, only $n=1$ works. Maybe $n=1$ is the only solution. … Selected Beam with score 0.998: First, let me understand what the problem is asking. It says that $n+2$ divides $3(n+3)(n^{2}+9)$. In other words, when we divide the product $3(n+3)(n^{2}+9)$ by $n+2$, the result should be an integer. So, I need to find all positive integers $n$ for which this condition holds and then sum them up. Let me write this out mathematically. If $n+2$ divides $3(n+3)(n^{2}+9)$, then there exists some integer $k$ such that: $3(n+3)(n^{2}+9)=k(n+2)$ But since we’re looking for integer solutions, maybe it’s better to think about this in terms of divisibility. So, $n+2$ must be a divisor of $3(n+3)(n^{2}+9)$. That suggests that $n+2$ divides this product. I remember that if a number $a$ divides another number $b$, then $b$ modulo $a$ is zero. So, maybe I can use modular arithmetic here. Let me see. Since $n+2$ divides $3(n+3)(n^{2}+9)$, it must be that: $3(n+3)(n^{2}+9)\equiv 0\pmod{n+2}$First, $n+3$ modulo $n+2$. Let’s substitute $n=-2+k$, but maybe that’s not necessary. Alternatively, we can note that $n+3=(n+2)+1$. So, modulo $n+2$, $n+3\equiv 1$. Similarly, $n^{2}+9$. Let me compute $n^{2}$ modulo $n+2$. Since $n\equiv-2\pmod{n+2}$, then $n^{2}\equiv(-2)^{2}=4\pmod{n+2}$. Therefore, $n^{2}+9\equiv 4+9=13\pmod{n+2}$. So, putting it all together, modulo $n+2$, the product becomes: $3(n+3)(n^{2}+9)\equiv 3\times 1\times 13=39\pmod{n+2}$ So, this implies that $39\equiv 0\pmod{n+2}$, which means that $n+2$ divides 39. Therefore, $n+2$ must be a positive divisor of 39. Now, the positive divisors of 39 are 1, 3, 13, and 39. But since $n$ is a positive integer, $n+2$ must be at least 3 (because if $n=1$, $n+2=3$). So, the possible values for $n+2$ are 3, 13, and 39. Therefore, the corresponding values for $n$ are: - If $n+2=3$, then $n=1$. - If $n+2=13$, then $n=11$. - If $n+2=39$, then $n=37$. …Therefore, the sum of all such positive integers $n$ is $1+11+37$. Let me compute that: 1 + 11 is 12, 12 + 37 is 49. …
@@
HMMT-25 Problem: Mark writes the expression $\sqrt{\underline{abcd}}$ on the board, where $\underline{abcd}$ is a four-digit number and $a\neq 0$. Derek, a toddler, decides to move the $a$, changing Mark’s expression to $a\sqrt{\underline{bcd}}$. Surprisingly, these two expressions are equal. Compute the only possible four-digit number $\underline{abcd}$. Rejected beam with score 0.337: Okay, so I have this problem where Mark writes the square root of a four-digit number, which is written as $\sqrt{(abcd)}$, where a isn’t zero. Then Derek moves the first digit ’a’ to the end, making the expression $a\sqrt{(bcd)}$. And the interesting part is that both expressions are equal. I need to find the only possible four-digit number $abcd$ that satisfies this condition. First, let me try to understand the problem step by step. Mark’s expression is the square root of a four-digit number, so that’s $\sqrt{(1000a+100b+10c+d)}$. Derek changes this by moving the first digit ’a’ to the end, so the new number is $1000b+100c+10d+a$, and Derek’s expression is a multiplied by the square root of this new number, which is $a\sqrt{(1000b+100c+10d+a)}$. And both of these are equal. So, mathematically, I can write this as: $\sqrt{(1000a+100b+10c+d)}=a\sqrt{(1000b+100c+10d+a)}$ If I square both sides to eliminate the square roots, I get: So, $1000a+100b+10c+d=a²*(1000b+100c+10d+a)$ Selected Beam with score 0.999: Let me write this down step by step. Let’s denote the four-digit number as $N=1000a+100b+10c+d$. So, Mark’s expression is $\sqrt{N}$. Then, Derek changes it to $a\sqrt{(bcd)}$, which is $a\sqrt{(100b+10c+d)}$. According to the problem, these two expressions are equal, Which means: $\sqrt{(1000a+100b+10c+d)}=a\sqrt{(100b+10c+d)}$ Let me square both sides to eliminate the square roots: 1000a + 100b + 10c + d = a² * (100b + 10c + d) Hmm, that gives me an equation: 1000a + 100b + 10c + d = a² (100b + 10c + d) Let me denote the four-digit number as N, so N = 1000a + 100b + 10c + d, and the three-digit number as M = 100b + 10c + d. So, the equation becomes: N = a² * M But since N is a four-digit number and a is a single digit (from 1 to 9, since $a\neq 0$), and M is a three-digit number (from 0 to 999). So, let me rewrite N as 1000a + M. So, 1000a + M = a² * M Let me rearrange this equation: 1000a = a² * M - M Factor out M on the right side: 1000a = M(a² - 1) Therefore, M = (1000a) / (a² - 1) Since M must be an integer (as it’s a three-digit number), (1000a) must be divisible by (a² - 1). So, (a² - 1) must be a divisor of 1000a. Let me note that a is a digit from 1 to 9. So, I can try each a from 1 to 9 and see if (a² - 1) divides 1000a, and then check if M is a three-digit integer. Let’s compute for each a: Starting with a=1: a=1: M = 1000*1 / (1 - 1) = 1000 / 0, which is undefined. So, a=1 is invalid. a=2: M = 1000*2 / (4 - 1) = 2000 / 3 $\approx$ 666.666…, which is not an integer. So, a=2 is invalid. a=3: M = 1000*3 / (9 - 1) = 3000 / 8 = 375. That’s an integer. So, M=375. a=4: M = 1000*4 / (16 - 1) = 4000 / 15 $\approx$ 266.666…, not integer. So, a=4 invalid. a=5: M = 1000*5 / (25 - 1) = 5000 / 24 $\approx$ 208.333…, not integer. a=6: M = 1000*6 / (36 - 1) = 6000 / 35 $\approx$ 171.428…, not integer. a=7: M = 1000*7 / (49 - 1) = 7000 / 48 $\approx$ 145.833…, not integer. a=8: M = 1000*8 / (64 - 1) = 8000 / 63 $\approx$ 127.450…, not integer. a=9: M = 1000*9 / (81 - 1) = 9000 / 80 = 112.5, which is not integer. So, only a=3 gives us an integer M=375, which is a three-digit number. Therefore, the only possible four-digit number is 3375. …

### v2

@@ Appendix D Further Details of Data Collection
Roll-in roll-out process. We also provide further intuition for the roll-in vs. roll-out process (illustrated in Fig.˜2 left). The roll-in and then roll-out process is a standard technique in imitation learning [12] and reinforcement learning [11].
@@
Roll-in. The main point of the roll-in process is to create a diverse distribution of partial solutions to roll-out from. By creating a diverse roll-in distribution with multiple roll-in policies, we can ensure that the classifier is trained on a diverse context distribution and will generalize better to new traces. To select where to cut a roll-in (to start the roll-out), we sample a cut index from the distribution of $p(i)=\frac{\sqrt{i}}{\sum_{j\in[L]}\sqrt{j}}$ where $L$ is the length of the roll-in. We chose this such that the cut index is more likely to occur at earlier positions of the roll-in. We want to encourage more learning at earlier positions since those prediction problems are more difficult than at later positions. The following figure (Fig.˜16) illustrates the distribution of the length of roll-outs, which shows that this indexing scheme indeed yields many long roll-outs.
@@ Appendix E Further Details for Value Model Training
Score Head Labels 0=Incorrect, 11=Correct, 22=Incomplete
@@
Grad Norm Clip 55
Compute GPUs $16$ nodes of $8\times\texttt{NVIDIA H100}$
@@ Appendix G Further Details for Training Bradley-Terry Reward Model
Dataset. Recall that our dataset for value model training (OpenR1-VM) contains $56$ responses per problem. To construct a Bradley-Terry dataset, we sample up to $4$ response pairs per problem, where a pair consists of a response with reward 0 (the ‘reject’ response) and a response with reward $1$ (the ‘chosen’ response). Some prompts may have fewer than $4$ responses with reward 0 or $1$, and in those cases, we include as many as possible. This yields a dataset of roughly 122122k pairs.
Model. We use the same model architecture as the value model (Appendix˜E) except that the score head outputs a single scalar score instead of a three-dimensional vector. We use the same training pipeline, except that the training loss is swapped to the standard BT loss [8]:
@@ Appendix H Inference FLOPS Computation
AIME-24 Problem: Jen enters a lottery by picking 44 distinct numbers from $S=\{1,2,3,\cdots,9,10\}.$ 44 numbers are randomly chosen from $S.$ She wins a prize if at least two of her numbers were 22 of the randomly chosen numbers, and wins the grand prize if all four of her numbers were the randomly chosen numbers. The probability of her winning the grand prize given that she won a prize is $\tfrac{m}{n}$ where $m$ and $n$ are relatively prime positive integers. Find $m+n$. Rejected beam with score 0.009: Okay, so I need to figure out the probability that Jen wins the grand prize given that she won a prize. That means I have to use conditional probability here. I remember that conditional probability is $P(A|B)=P(A\cap B)/P(B)$. In this case, event A is her winning the grand prize, and event B is her winning a prize. So, I need to find the probability that she won both the grand prize and a prize, which is just her winning the grand prize, divided by the probability that she won a prize. First, let me figure out the total number of possible ways Jen can pick 4 numbers from S, which has 10 numbers. That should be the combination of 10 choose 4, right? So, $C(10,4)=210$. So, the total number of possible combinations is 210. So, she can win a prize in two scenarios: either exactly two of her numbers are the randomly chosen ones, or all four are. …The number of ways she can have exactly two numbers correct is $C(4,2)$ ways to choose the correct numbers, and then $C(6,2)$ ways to choose the incorrect ones from the remaining 6 numbers. Similarly, the number of ways she can have all four numbers correct is just $C(4,4)=1$. So, the total number of winning ways is C(4,2)C(6,2)+C(4,4)C(4,2)*C(6,2)+C(4,4).C(4,2)C(4,2) is 6, and $C(6,2)$ is 15. So, 6*15 = 90. Then, $C(4,4)$ is 1. So, total winning ways are 90 + 1 = 91. Alternatively, I can compute the probability of not winning a prize, which is the probability that she has 0 or 1 correct numbers. Let me compute that and subtract from 1 to see if it matches. Number of ways to have 0 correct numbers: C(4,0)C(6,4)=115=15C(4,0)*C(6,4)=1*15=15. Number of ways to have 1 correct number: C(4,1)C(6,3)=420=80C(4,1)*C(6,3)=4*20=80. So, total non-winning ways: 15 + 80 = 95. Therefore, probability of not winning a prize is 95/210, so probability of winning a prize is 195/210=(21095)/210=115/2101-95/210=(210-95)/210=115/210. Wait, that doesn’t match with the previous result. Hmm, something’s wrong here. Wait, why did I get 91 before? Let me see. I had C(4,2)C(6,2)+C(4,4)C(4,2)*C(6,2)+C(4,4). So, that’s 6*15 +1=91. But that’s only the number of winning ways where she has exactly two or exactly four correct numbers. So, 91 winning ways, 95 non-winning, which adds to 186, which is less than 210. So, the rest must be the impossible case? No, because in reality, she can’t have 0 correct numbers if she has 4 numbers. Wait, no, she can have 0 correct numbers if all her numbers are among the 6 incorrect ones. So, that’s 15 ways. Similarly, 1 correct number is 80 ways, and 2 correct numbers is 90 ways, and 4 correct numbers is 1 way. So, 15 +80 +90 +1=186. Wait, so actually, 186 winning ways and 24 non-winning ways. So, probability of winning is 186/210, which simplifies to 31/35. Therefore, $P(A|B)=(1/210)/(186/210)=1/186$. So, m is 1 and n is 186. Are they relatively prime? 1 and 186 share no common factors besides 1, so yes. So, $m+n=1+186=187$. Selected Beam with score 0.996: Okay, so this is a conditional probability problem. The formula for conditional probability is $P(A|B)=P(A\cap B)/P(B)$, where A is the event that she wins the grand prize, and B is the event that she wins a prize. Since winning the grand prize is a specific case of winning a prize, A is a subset of B. So, $P(A\cap B)$ is just $P(A)$. Therefore, the probability we’re looking for is $P(A)/P(B)$. Since she picks 4 numbers from 10, the total number of possible 4-number combinations is $C(10,4)$. The number of favorable outcomes for A is 1 (since there’s only one exact set of four numbers that could be the randomly chosen ones). So, $P(A)$ is $1/C(10,4)$. To find $P(B)$, it might be easier to compute the probability of the complementary event, which is that Jen wins zero prizes or one prize, and subtract that from 1. So, $P(B)$ = 1 - P(0 prizes) - P(1 prize). Let me compute P(0 prizes) and P(1 prize). First, P(0 prizes): This is the probability that none of Jen’s four numbers are among the four randomly chosen. So, all four numbers she picked are from the remaining six numbers (since S has 10 numbers, and if she picks 4, the remaining are 6). So, the number of ways to choose 4 numbers from 6 is $C(6,4)$. Therefore, P(0 prizes) = $C(6,4)/C(10,4)$. Similarly, P(1 prize): This is the probability that exactly one of Jen’s numbers is among the four randomly chosen. So, we need to choose 1 number from her 4, and 3 numbers from the remaining 6 (since 10 - 4 = 6). So, the number of ways is C(4,1)C(6,3)C(4,1)*C(6,3). Therefore, P(1 prize) = [C(4,1)C(6,3)]/C(10,4)[C(4,1)*C(6,3)]/C(10,4). …So, putting it all together: P(B)=1[C(6,4)/C(10,4)][C(4,1)C(6,3)/C(10,4)]P(B)=1-[C(6,4)/C(10,4)]-[C(4,1)*C(6,3)/C(10,4)] Therefore, P(B)=115/21080/210P(B)=1-15/210-80/210. …Therefore, $P(B)$ is 23/42. …Therefore, the conditional probability $P(A|B)$ is (1/210)/(23/42)=(1/210)(42/23)=(42)/(21023)(1/210)/(23/42)=(1/210)*(42/23)=(42)/(210*23) …Therefore, the probability becomes (1/5)/23=1/(523)=1/115(1/5)/23=1/(5*23)=1/115. …But 1/115 can’t be simplified further, so m is 1 and n is 115, which are coprime. Therefore, m + n is 1 + 115 = 116.
@@
AIME-25 Problem: Find the sum of all positive integers $n$ such that $n+2$ divides the product $3(n+3)(n^{2}+9)$. Rejected beam with score 0.395: First, let’s parse the problem. I need to find all positive integers $n$ where $n+2$ is a divisor of $3(n+3)(n^{2}+9)$. So, mathematically, this means that $n+2$ divides $3(n+3)(n^{2}+9)$. In other words, when you divide $3(n+3)(n^{2}+9)$ by $n+2$, there should be no remainder. So, $n+2$ must be a factor of this product. One way to approach this is by using the Factor Theorem. The Factor Theorem states that if $n+2$ is a factor of a polynomial $P(n)$, then P(2)=0P(-2)=0. So, if I let $P(n)=3(n+3)(n^{2}+9)$, then P(2)P(-2) should be equal to zero if $n+2$ divides $P(n)$. Let me compute P(2)P(-2): $P(-2)=3(-2+3)((-2)^{2}+9)$ Simplify each part: (2+3)=1(-2+3)=1 $(-2)^{2}=4$, so $4+9=13$ Thus, P(2)=3113=39P(-2)=3*1*13=39Wait, perhaps I’m misapplying the Factor Theorem here. Let me think again. Alternatively, another approach is to express $n+2$ as $k$, so $k=n+2$. Then, n=k2n=k-2. Substitute n=k2n=k-2 into the expression $3(n+3)(n^{2}+9)$ and see what conditions must hold for $k$ to divide this. …Therefore, the product becomes: $3(k+1)(k^{2}-4k+13)$ Which implies that $k$ divides $3(k+1)(k^{2}-4k+13)$. Since $k$ divides the product, it must divide at least one of the factors. But $k$ divides $k$, so we can consider the other factors. But wait, $k$ divides $3(k+1)(k^{2}-4k+13)$, so $k$ divides $3(k+1)(k^{2}-4k+13)$. Since $k$ and $k+1$ are consecutive integers, their GCD is 1. Similarly, $k$ and $k^{2}-4k+13$ may have a GCD. Let me check: Let me denote $d=\gcd(k,k^{2}-4k+13)$. Then $d$ divides $k$ and $k^{2}-4k+13$. So, $d$ divides $k^{2}-4k+13-k(k-4)=k^{2}-4k+13-k^{2}+4k=13$. So, $d$ divides 13. Therefore, $d=1$ or $13$. Wait, maybe I’m complicating things. Let me approach this differently. … But maybe there’s another way to approach this. Let me try plugging in small positive integers for $n$ and see if $n+2$ divides the product $3(n+3)(n^{2}+9)$. That might help confirm if $n=1$ is the only solution. Testing $n=1$: As above, product is 120, which is divisible by 3. Testing $n=2$: $n+2=4$. Product is $3(5)(13)=195$. 195 divided by 4 is 48.75, which isn’t an integer. So, 4 doesn’t divide 195. Testing $n=9$: $n+2=11$. Product is $3(12)(90)=3240$. 3240 divided by 11 is approximately 294.545, not integer. Testing $n=10$: $n+2=12$. Product is $3(13)(109)=4209$. 4209 divided by 12 is 350.75, not integer. Wait, so up to $n=10$, only $n=1$ works. Maybe $n=1$ is the only solution. Selected Beam with score 0.998: First, let me understand what the problem is asking. It says that $n+2$ divides $3(n+3)(n^{2}+9)$. In other words, when we divide the product $3(n+3)(n^{2}+9)$ by $n+2$, the result should be an integer. So, I need to find all positive integers $n$ for which this condition holds and then sum them up. Let me write this out mathematically. If $n+2$ divides $3(n+3)(n^{2}+9)$, then there exists some integer $k$ such that: $3(n+3)(n^{2}+9)=k(n+2)$ But since we’re looking for integer solutions, maybe it’s better to think about this in terms of divisibility. So, $n+2$ must be a divisor of $3(n+3)(n^{2}+9)$. That suggests that $n+2$ divides this product. I remember that if a number $a$ divides another number $b$, then $b$ modulo $a$ is zero. So, maybe I can use modular arithmetic here. Let me see. Since $n+2$ divides $3(n+3)(n^{2}+9)$, it must be that: $3(n+3)(n^{2}+9)\equiv 0\pmod{n+2}$ …First, $n+3$ modulo $n+2$. Let’s substitute n=2+kn=-2+k, but maybe that’s not necessary. Alternatively, we can note that $n+3=(n+2)+1$. So, modulo $n+2$, $n+3\equiv 1$. Similarly, $n^{2}+9$. Let me compute $n^{2}$ modulo $n+2$. Since $n\equiv-2\pmod{n+2}$, then $n^{2}\equiv(-2)^{2}=4\pmod{n+2}$. Therefore, $n^{2}+9\equiv 4+9=13\pmod{n+2}$. So, putting it all together, modulo $n+2$, the product becomes: $3(n+3)(n^{2}+9)\equiv 3\times 1\times 13=39\pmod{n+2}$ So, this implies that $39\equiv 0\pmod{n+2}$, which means that $n+2$ divides 39. Therefore, $n+2$ must be a positive divisor of 39. Now, the positive divisors of 39 are 1, 3, 13, and 39. But since $n$ is a positive integer, $n+2$ must be at least 3 (because if $n=1$, $n+2=3$). So, the possible values for $n+2$ are 3, 13, and 39. Therefore, the corresponding values for $n$ are: - If $n+2=3$, then $n=1$. - If $n+2=13$, then $n=11$. - If $n+2=39$, then $n=37$. …Therefore, the sum of all such positive integers $n$ is $1+11+37$. Let me compute that: 1 + 11 is 12, 12 + 37 is 49.
@@
HMMT-25 Problem: Mark writes the expression $\sqrt{\underline{abcd}}$ on the board, where $\underline{abcd}$ is a four-digit number and $a\neq 0$. Derek, a toddler, decides to move the $a$, changing Mark’s expression to $a\sqrt{\underline{bcd}}$. Surprisingly, these two expressions are equal. Compute the only possible four-digit number $\underline{abcd}$. Rejected beam with score 0.337: Okay, so I have this problem where Mark writes the square root of a four-digit number, which is written as $\sqrt{(abcd)}$, where a isn’t zero. Then Derek moves the first digit ’a’ to the end, making the expression $a\sqrt{(bcd)}$. And the interesting part is that both expressions are equal. I need to find the only possible four-digit number abcdabcd that satisfies this condition. First, let me try to understand the problem step by step. Mark’s expression is the square root of a four-digit number, so that’s $\sqrt{(1000a+100b+10c+d)}$. Derek changes this by moving the first digit ’a’ to the end, so the new number is $1000b+100c+10d+a$, and Derek’s expression is a multiplied by the square root of this new number, which is $a\sqrt{(1000b+100c+10d+a)}$. And both of these are equal. So, mathematically, I can write this as: $\sqrt{(1000a+100b+10c+d)}=a\sqrt{(1000b+100c+10d+a)}$ If I square both sides to eliminate the square roots, I get: So, 1000a+100b+10c+d=a²(1000b+100c+10d+a)1000a+100b+10c+d=a²*(1000b+100c+10d+a) Selected Beam with score 0.999: Let me write this down step by step. Let’s denote the four-digit number as $N=1000a+100b+10c+d$. So, Mark’s expression is $\sqrt{N}$. Then, Derek changes it to $a\sqrt{(bcd)}$, which is $a\sqrt{(100b+10c+d)}$. According to the problem, these two expressions are equal, Which means: $\sqrt{(1000a+100b+10c+d)}=a\sqrt{(100b+10c+d)}$ Let me square both sides to eliminate the square roots: 1000a + 100b + 10c + d = a² * (100b + 10c + d) Hmm, that gives me an equation: 1000a + 100b + 10c + d = a² (100b + 10c + d) Let me denote the four-digit number as N, so N = 1000a + 100b + 10c + d, and the three-digit number as M = 100b + 10c + d. So, the equation becomes: N = a² * M But since N is a four-digit number and a is a single digit (from 1 to 9, since $a\neq 0$), and M is a three-digit number (from 0 to 999). So, let me rewrite N as 1000a + M. So, 1000a + M = a² * M Let me rearrange this equation: 1000a = a² * M - M Factor out M on the right side: 1000a = M(a² - 1) Therefore, M = (1000a) / (a² - 1) Since M must be an integer (as it’s a three-digit number), (1000a) must be divisible by (a² - 1). So, (a² - 1) must be a divisor of 1000a. Let me note that a is a digit from 1 to 9. So, I can try each a from 1 to 9 and see if (a² - 1) divides 1000a, and then check if M is a three-digit integer. Let’s compute for each a: Starting with a=1: a=1: M = 1000*1 / (1 - 1) = 1000 / 0, which is undefined. So, a=1 is invalid. a=2: M = 1000*2 / (4 - 1) = 2000 / 3 $\approx$ 666.666…, which is not an integer. So, a=2 is invalid. a=3: M = 1000*3 / (9 - 1) = 3000 / 8 = 375. That’s an integer. So, M=375. a=4: M = 1000*4 / (16 - 1) = 4000 / 15 $\approx$ 266.666…, not integer. So, a=4 invalid. a=5: M = 1000*5 / (25 - 1) = 5000 / 24 $\approx$ 208.333…, not integer. a=6: M = 1000*6 / (36 - 1) = 6000 / 35 $\approx$ 171.428…, not integer. a=7: M = 1000*7 / (49 - 1) = 7000 / 48 $\approx$ 145.833…, not integer. a=8: M = 1000*8 / (64 - 1) = 8000 / 63 $\approx$ 127.450…, not integer. a=9: M = 1000*9 / (81 - 1) = 9000 / 80 = 112.5, which is not integer. So, only a=3 gives us an integer M=375, which is a three-digit number. Therefore, the only possible four-digit number is 3375.